OpenStudy (amtran_bus):

Solve by elimination: The perimeter of a rectangular floor is 90 feet. Find the dimensions of the floor if the length is twice the width.

OpenStudy (amtran_bus):

How would you even set it up?

OpenStudy (anonymous):

are you still there amtran?

OpenStudy (amtran_bus):

Yes I am.

OpenStudy (anonymous):

word problems go better one step at a time. So, first step. what is the formula for the perimeter of a rectangle (because this is a rectangular floor).

OpenStudy (amtran_bus):

2L+2w

OpenStudy (anonymous):

right... nice

OpenStudy (anonymous):

now, how can we make an equation that states that with the given info in the first sentence?

OpenStudy (amtran_bus):

2l+2w=90?

OpenStudy (anonymous):

right, but notice how there is no involvement of area in this problem

OpenStudy (anonymous):

yes... 2L+2W=90 is perfect!!

OpenStudy (anonymous):

what about the second sentence?

OpenStudy (amtran_bus):

But using elimination, we have to have another equation. Would it be 2l+w=90?

OpenStudy (amtran_bus):

That says 2L not 21 sorry

OpenStudy (anonymous):

i understand.. lol that is the way i read it.

OpenStudy (anonymous):

Read the second sentence and see if you can come up with another equation to make elimination possible.

OpenStudy (anonymous):

brb

OpenStudy (amtran_bus):

2l+2w=90 2L+w=180? I have a hard time setting them up!

OpenStudy (anonymous):

it says in the second sentence that the length is twice the width. how would you write that in equation form?

OpenStudy (anonymous):

remember that is means equals

OpenStudy (anonymous):

if the length is two times the width, that means that L= ?

jhonyy9 (jhonyy9):

why is this so like difficile ? you know that P=2L +2W and that 1L=2W than this is very easy

OpenStudy (anonymous):

i was trying to help him get to the answer, not give it to him.

jhonyy9 (jhonyy9):

yes ok i understand it so step by step will be very very good .correct

OpenStudy (anonymous):

so now amtram, you can see that jhonyy set up the system.

OpenStudy (anonymous):

it doesnt matter to me. i just want you to understand how to do this problem and others like it. give it to jhonyy

OpenStudy (amtran_bus):

So you set it up just like that? And solve? Will you plug it all in and see what it comes to be? I have 30 and 15.

OpenStudy (anonymous):

that would be correct

OpenStudy (amtran_bus):

Will you draw it out for me?

OpenStudy (anonymous):

L=30 and W=15

OpenStudy (anonymous):

do you want me to make a drawing or give you a set of instructions?

OpenStudy (amtran_bus):

using the draw feature please.

OpenStudy (anonymous):

unfortunately, the draw feature is not working properly for me. but i can give you instructions on how to draw your own picture if you like

OpenStudy (amtran_bus):

thanks

OpenStudy (anonymous):

draw a rectangle. it does not have to be perfect.

OpenStudy (amtran_bus):

k

OpenStudy (anonymous):

lol.

OpenStudy (anonymous):

now label the short sides w and the long sides l.

OpenStudy (amtran_bus):

k

OpenStudy (anonymous):

then on the long sides, set the l=2w

OpenStudy (amtran_bus):

k

OpenStudy (anonymous):

lol... so then, you add up the lengths of the sides to get the perimeter length and set those lengths equal to 90. back to 2L + 2W=90

OpenStudy (anonymous):

this problem is more suited to substitution because L=2w, but you can use elimination here too.

OpenStudy (amtran_bus):

Thanks!

OpenStudy (anonymous):

by subtracting 2w from both sides, you get L-2w=0

OpenStudy (anonymous):

now you have L-2w=0 and 2L+2w=90

OpenStudy (anonymous):

then you can add the two equations together to get 3L=90 or L=30

OpenStudy (amtran_bus):

but how do u get the actual #'s?

OpenStudy (anonymous):

when you add two numbers together, you put one number above the other and line them up, right?

OpenStudy (amtran_bus):

yes